# Further author information: (Send correspondence to Yannik Mahlau)

### v1

@@ 3.1 Speed Comparison
In Tab. 2, the runtime of different simulation frameworks is displayed for this setup. This comparison is highly dependent on the hardware used and can change drastically. We tried to make this comparison as fair as possible by using the strongest hardware available in our research group, which is compatible with the respective frameworks. The open-source packages Ceviche and Meep were run on a CPU-Cluster with 16 cores (AMD EPYC, 4137MHz) and 256GB RAM. For the highest resolution, both frameworks reached the memory limitations of this system. OmniSim was run on a Windows consumer computer (Intel i7-8550U CPU/1.80 GHz, 16 GB RAM) due to licensing. Resulting from the low memory, OmniSim was only able to run the simulations up to a resolution of $20nm$. All three frameworks would be able to simulate finer resolutions on stronger hardware. In contrast to the previous frameworks, Tidy3D uses a web-based service, where users can submit their simulation jobs online. These jobs are run on undisclosed special-purpose hardware, which makes comparison difficult. Additionally, runtimes may vary depending on the current demand. For example, for the resolution of $10nm$, we observed a runtime between 1 minute 30 seconds and 2 minutes. The values reported for Tidy3D in Tab. 2 are average values for 3 runs at different days and times of day for resolutions $25$, $20$, and $10nm$. For $5nm$, we performed only a single run for cost reasons. Our license is limited to 800 million grid points, such that we were not able to test the simulation at the finest resolution. We would like to emphasize that the pay-per-simulation model of Tidy3D would result in high costs for research projects with large simulation demand. This makes running automatic optimization requiring multiple hundreds of simulations infeasible. In contrast, our open-source FDTDX software can be run on in-house GPU, university compute clusters, or compute services like AWS if necessary. However, using compute services will also result in costs which depend on the hardware used. We ran the FDTDX simulation on a single Nvidia H100 GPU for all resolutions up to $5nm$. The simulation with $2.5nm$ resolution was run on four Nvidia H100 GPUs due to larger memory requirements. The results of the speed comparison show that FDTDX outperforms all other open-source simulation software by a large margin. At 288 million grid cells, it achieves a speedup of about 10x compared to Meep and 415x compared to Ceviche. For 36 million grid cells, the speed of FDTDX is close to or even exceeds the speed of Tidy3D in some runs. Only in very large simulations is FDTDX significantly slower than Tidy3D. However, in practice, this speed difference only matters when running many simulations at this size, which would be cost-prohibitive using Tidy3D. Moreover, running very large simulations may not even be possible without special licenses.
@@
In this speed comparison, we did not include any setup times. For Tidy3D, we excluded the time required to send and setup the simulation and receive the results. Additionally, we did not include the waiting time in queue if demand is high. For FDTDX, we did not include the time required to compile and automatically optimize the computational graph in JAX. This compilation step took between a single second for 2.3 million grid cells and two minutes for 2.3 billion grid cells.

### v2

@@ 3.1 Speed Comparison
In Tab. 2, the runtime of different simulation frameworks is displayed for this setup. This comparison is highly dependent on the hardware used and can change drastically. We tried to make this comparison as fair as possible by using the strongest hardware available in our research group, which is compatible with the respective frameworks. The open-source packages Ceviche and Meep were run on a CPU-Cluster with 16 cores (AMD EPYC, 4137MHz) and 256GB RAM. For the highest resolution, both frameworks reached the memory limitations of this system. OmniSim was run on a Windows consumer computer (Intel i7-8550U CPU/1.80 GHz, 16 GB RAM) due to licensing. Resulting from the low memory, OmniSim was only able to run the simulations up to a resolution of $20nm$. All three frameworks would be able to simulate finer resolutions on stronger hardware. In contrast to the previous frameworks, Tidy3D uses a web-based service, where users can submit their simulation jobs online. These jobs run on undisclosed special-purpose hardware, making comparisons difficult. Additionally, runtimes may vary depending on the current demand. For example, for the resolution of $10nm$, we observed a runtime between 1 minute 30 seconds and 2 minutes. The values reported for Tidy3D in Tab. 2 are average values for 3 runs at different days and times of day for resolutions $25$, $20$, and $10nm$. For $5nm$, we performed only a single run for cost reasons. For the run at $2.5nm$, we generously received funding from the Tidy3D developers themselves as a standard license is limited to 800 million grid points. The pay-per-simulation model of Tidy3D would result in high costs for research projects with a large simulation demand. This is exacerbated for gradient-based optimization which require many forward and backward simulations. In contrast, our open-source FDTDX software can be run on in-house GPU, university compute clusters, or compute services like AWS if necessary. However, using compute services will also result in costs which depend on the hardware used. Since a cost comparison is difficult and highly dependent on the service provider, we do not consider it here. We ran the FDTDX simulation on a single Nvidia H100 GPU for all resolutions up to $5nm$. The simulation with $2.5nm$ resolution was run on four Nvidia H100 GPUs due to larger memory requirements. The results of the speed comparison show that FDTDX outperforms all other open-source simulation software by a large margin. At 288 million grid cells, it achieves a speedup of about 10x compared to Meep and 415x compared to Ceviche. For 36 million grid cells, the speed of FDTDX is close to or even exceeds the speed of Tidy3D in some runs. Only in very large simulations is FDTDX significantly slower than Tidy3D. However, without calculating exact values, it is reasonable to assume that for large simulations, the price difference increases with the speed difference. We believe that there is a place for both open-source and commercial software as open-source software is usually more cost-effective and gives researchers more freedom to implement their own algorithms. In contrast, commercial software is often faster, but with the development of FDTDX, we make a significant step towards decreasing this gap.
@@
In this speed comparison, we did not include any setup times. For Tidy3D, we excluded the time required to send and setup the simulation and receive the results. Additionally, we did not include waiting time in queue if demand is high. For both Tidy3D and FDTDX, we did not include the time required to compile and optimize the computational graph. This compilation step took between a single second for 2.3 million grid cells and two minutes for 2.3 billion grid cells.